HETAV PATEL

Grade 11,

How many time constants will elapse before the charge on a capacitor falls to 0.1% of its maximum value in a discharging RC circuit?

How many time constants will elapse before the charge on a capacitor falls to 0.1% of its maximum value in a discharging RC circuit?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. q = Qe^–t/RC q = 0.1 % Q RC ⇒ Time constant = 1 * 10^–3 Q So, 1 * 10^–3 Q = Q * e^–t/RC ⇒ e^–t/RC = ln 10^–3 ⇒ t/RC = –(–6.9) = 6.9